# Versatile Motion-Language Models for Multi-turn Interactive Agents

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have greatly enhanced their ability to generate natural and contextually relevant text, making AI interactions more human-like. However, generating and understanding interactive human-like motion, where two individuals engage in coordinated movements, remains a challenge due to the complexity of modeling these coordinated interactions. Furthermore, a versatile model is required to handle diverse interactive scenarios, such as chat systems that follow user instructions or adapt to their assigned role while adjusting interaction dynamics. To tackle this problem, we introduce VIM, short for the Versatile Interactive Motion language model, which integrates both language and motion modalities to effectively understand, generate, and control interactive motions in multi-turn conversational contexts. To address the scarcity of multi-turn interactive motion data, we introduce a synthetic dataset called Inter-$\text{MT}^{2}$; where we utilize pre-trained models to create diverse instructional datasets with interactive motion. Our approach first trains a motion tokenizer that encodes interactive motions into residual discrete tokens. In the pre-training stage, the model learns to align motion and text representations with these discrete tokens. During the instruction fine-tuning stage, VIM adapts to multi-turn conversations using Inter-$\text{MT}^{2}$. We evaluate the versatility of our method across motion-related tasks—motion-to-text, text-to-motion, reaction generation, motion editing, and reasoning about motion sequences. The results highlight VIM’s versatility and effectiveness in handling complex interactive motion synthesis.
@@ 1 Introduction
Agents that reflect how humans communicate and interact with each other through motion have the potential to revolutionize our interaction with technology. By capturing the subtleties of human communication, including gestures, expressions, and interactive behaviors, these agents can offer more intuitive and natural interfaces. This holistic understanding enables technology to adjust its responses and behaviors based on the user’s physical motions and situational context, leading to more personalized and engaging interactions. Such capabilities are crucial for enhancing support across various domains, including robotics, virtual humans, entertainment, and more.
@@
Recent advancements in large language models (LLMs) (Dubey et al., 2024; Team et al., 2024; Yang et al., 2024) have demonstrated significant potential in generating human-like text and understanding complex linguistic interactions. They have even extended their capability to multi-modal contexts, successfully integrating various input sources such as images, speech, and videos (Ge et al., 2024; Liu et al., 2024; Chen et al., 2024b; Tang et al., 2024; Shu et al., 2023). Building upon these developments, there is a growing interest in incorporating human (or robot) motion as a new modality (Jiang et al., 2024; Chen et al., 2024a), leading to the emergence of the “motion-language models” (MLM). However, existing approaches (Zhang et al., 2023; Guo et al., 2024a; 2022; Zhang et al., 2024d; Cai et al., 2024) often focus on single tasks, such as text-to-motion or motion-to-text translation, and consider only single motions without interactions. This limitation hinders the agents’ ability to handle scenarios involving multi-agents, complex interactions, and multi-turn conversations.
@@
Beyond modeling the motion of a single person, interactive motion, where two individuals participate in interactions, allows the model to learn about social behavior. Modeling such interactions requires versatility to effectively control interactions, allowing users to provide instructions, assign roles, or modify behaviors. In this paper, we aim to build a unified yet versatile motion-language model designed to generate, control, and comprehend sophisticated interactive motions.
@@
One of the primary challenges in developing these interactive agents is the lack of multi-turn interactive motion data. Datasets containing motions of two individuals interacting with each other, along with multi-turn conversational instructions, are scarce and challenging to collect. This makes it difficult for models to learn the nuances of interactive motions and multi-turn dynamics.
To address this, we present a new synthesized dataset called Inter-$\text{MT}^{2}$, which includes various instructions about the interactive motions, in a multi-turn conversational format. We utilize a large language model to produce diverse instructions with motion captions and a diffusion-based text-to-motion model to generate corresponding motions. We expect that leveraging such foundation models to construct training data allows the model to generalize more effectively with prior knowledge.
@@
Building upon our synthesized dataset, we present VIM, a Versatile Interactive Motion-language model designed for multi-turn conversations involving interactive motions. We pursue the versatility of VIM through a unified architecture that can simultaneously input and output both motion and text modalities. Based on the pre-trained LLMs, our training process can be divided into three stages: (1) training of the interaction motion tokenizer, (2) pre-training for motion and text representation alignment, and (3) instruction tuning with our synthesized dataset, Inter-$\text{MT}^{2}$, to handle more complex and multi-turn instructions. This enables VIM to effectively comprehend, generate, and control interactive motions, as illustrated in Figure 1. To assess VIM’s capabilities, we introduce new evaluation protocols that evaluate its performance across various motion-related tasks. This include editing motions and reasoning about motion sequences based on contextual cues, highlighting its versatility and effectiveness in complex motion interaction scenarios.
Figure 1: We introduce VIM, a versatile interactive motion language model. Left: reaction generation, motion reasoning, and generation. Right top: text-to-motion generation. Right bottom: Motion editing and motion understanding.
@@
In summary, the main contributions of this paper are threefold: (1) We propose VIM that can simultaneously process and generate both motion and text modalities, along with a three-stage training pipeline consisting of motion tokenizer training, pre-training for modality alignment, and instruction tuning. (2) We present Inter-$\text{MT}^{2}$, a multi-turn interactive motion-text dataset, to address the lack of multi-turn interactive motion data. (3) We introduced a new evaluation protocol to evaluate the performance of motion-language models on complex motion interaction scenarios.
@@ 2 Related Work
Advancements in human motion modeling have driven significant progress in both motion generation and control. Diffusion-based methods, such as MDM (Tevet et al., 2023), FG-T2M (Wang et al., 2023), and MotionDiffuse (Zhang et al., 2024a) excel at synthesizing realistic human motions from the text. Transformer-based models with vector quantization, such as TM2T (Guo et al., 2022) and T2M-GPT (Zhang et al., 2023), effectively capture complex motion patterns. MoMASK (Guo et al., 2024a) improves motion granularity with residual tokenizers. For motion editing, some approaches focus on style transfer (Aberman et al., 2020; Guo et al., 2024b) or specific body part modifications (Zhang et al., 2024a; Kim et al., 2023). FineMoGEN (Zhang et al., 2024c) offers fine-grained motion synthesis based on user instructions. MEOs (Goel et al. (2024)) use captions and large language models to identify frames and body parts to edit, while MotionFix (Athanasiou et al. (2024)) conditions diffusion models on both source motion and edit text for seamless motion edits. However, these models usually target single tasks (e.g., text-to-motion, or motion editing) and lack versatility in handling input and output of both motion and text simultaneously in a unified architecture.
Motion Language Model
Recent developments in motion language models have aimed to achieve versatility across various motion-related tasks. MotionGPT (Jiang et al., 2023) demonstrates versatility in motion comprehension and generation based on a unified framework. MotionChain (Jiang et al., 2024) introduces a multi-turn conversational system for interpreting and generating motions within dialogue contexts, including image inputs. Zhou et al. (2024) introduces AvatarGPT integrating motion generation and planning ability in motion large language model. Some studies, like Chen et al. (2024a), expand modalities into speech, music, and videos but focus primarily on comprehension rather than generation. Zhang et al. (2024b) proposed unified models for generating motion from various input modalities. M3-GPT, from Luo et al. (2024), models speech, music, text, and motion interchangeably. However, modeling interactive motions in versatile large models remains under-explored. While some efforts, such as Wu et al. (2024), address this direction, they often lack multi-turn interactions and complex reasoning abilities. Our work addresses this gap with a model trained on our synthesized Inter-$\text{MT}^{2}$ dataset, enabling the understanding and generation of interactive motions in multi-turn conversations with advanced reasoning capabilities. This approach facilitates more nuanced, context-aware motion generation in complex interactive behaviors.
@@
Modeling human-human interactions has garnered increasing attention in recent research. Several multi-person interaction datasets (Ng et al., 2020; Fieraru et al., 2020; Yin et al., 2023) have been developed, and recent efforts like Inter-X (Xu et al., 2024a) and InterHuman (Liang et al., 2024) have collected interactive motions paired with textual descriptions for text-based motion control. In text-to-motion tasks, InterGEN (Xu et al., 2024a) introduces a diffusion-based model with spatial constraint loss. PriorMDM (Shafir et al., 2024) leverages pre-trained motion diffusion models with slim communication blocks. For reaction generation, ReMoS Ghosh et al. (2023) synthesizes reactive motion using spatio-temporal cross-attention, while ReGenNet Xu et al. (2024b) employs a transformer-based model with distance-based interaction loss to predict human reactions. While existing models have advanced interactive motion modeling, they lack versatility and focus on specific tasks, failing to capture complex multi-turn dynamics. To address this, we introduce Inter-$\text{MT}^{2}$, enabling agents to generate sophisticated motions, respond to instructions, adapt roles, and adjust behaviors based on context.
@@ 3 $\text{Inter}\text{-MT}^{2}$: Interactive multi-turn motion-text dataset
Current datasets (Yin et al., 2023; Liang et al., 2024; Xu et al., 2024a) for modeling interactive motions lack sufficient diversity in instructions and do not include multi-turn conversations. To address this gap, we introduce Inter-$\text{MT}^{2}$: interactive muti-turn motion-text dataset. This dataset covers a variety of interactive motion scenarios with multi-turn conversations, diverse instructions, and spatiotemporally aligned motions between two individuals.
Figure 2: Overview of synthetic data generation for multi-turn conversations with interactive motions. (a) Motion captions and instructions are generated using GPT-4 based on interactions between two characters, followed by (b) the corresponding motion being synthesized using the InterGEN.
Collecting diverse instructional data for interactive motions and multi-turn conversation samples poses significant challenges. The first challenge is obtaining instruction and conversational data that encompass complex reasoning and generation capabilities. Leveraging pre-trained foundation models to generate a broad range of instructions can enrich the dataset with diverse and intricate examples. The second challenge is acquiring motion data that aligns with these text instructions. Functional approaches using rule-based methods may struggle to maintain spatial and temporal constraints in complex interactive scenarios, while retrieval-based methods are limited by dependence on existing cases and lack diversity. Alternatively, generative approaches using pre-trained models show promise in producing diverse, complex text-to-motion sequences, offering more flexibility for modeling interactive motions.
We utilize the Inter-X (Xu et al., 2024a) and InterHuman (Liang et al., 2024) datasets as the foundation for building our datasets. We further employ GPT-4o (OpenAI, 2024) to generate motion captions and conversational instructions for a variety of tasks, such as motion editing, reasoning, and story generation, enhancing the model’s versatility. Motion captions are sourced from these base datasets or generated by large language models (LLMs). We utilize the state-of-the-art text-to-motion diffusion model, InterGEN (Liang et al., 2024), to generate corresponding motions that align with the generated caption from LLMs. Our data collection pipeline, shown in Figure 2, comprises 82K samples of multi-turn conversational data involving interactive motions, including 96K of synthesized interactive motions and 56K motions from the source dataset.
@@ 4 VIM: Versatile Interactive Motion-Language Model
In this section, we introduce VIM, a versatile interactive motion language model, designed to incorporate multi-turn conversations considering both language and interactive motion as input or output modality. First, we will explain the underlying philosophy behind our design choices for the model architectures, followed by a detailed description of the training methodologies. Then, we introduce advanced interactive motion tasks in multi-turn conversations.
@@ 4.1 Notations
Formally, we denote interactive motion from two individual $a$ and $b$ as $\{\mathbf{m}_{a},\mathbf{m}_{b}\}$, following non-canonical representation from Liang et al. (2024) based on SMPL-X structure (Pavlakos et al., 2019). Each timestep of the motion $\mathbf{m}^{i}=[\mathbf{j}_{g}^{p},\mathbf{j}_{g}^{v},\mathbf{j}^{r},\mathbf{c}^{f}]$ is composed of global joint positions $\mathbf{j}_{g}^{p}\in\mathbb{R}^{3N_{j}}$, global joint velocities $\mathbf{j}_{g}^{v}\in\mathbb{R}^{3N_{j}}$, 6D representation of local rotations $\mathbf{j}^{r}\in\mathbb{R}^{6N_{j}}$, with the number of joints $N_{j}$, and binary ground contact features $\mathbf{c}^{f}\in\mathbb{R}^{4}$. We aim to train a motion language model $p_{\theta}$ which can model texts and motions for both inputs and outputs. We define input as an instruction or previous context and output as an answer, with the template below
[…]
where $X_{u}$ and $X_{a}$ are composed of both a mixture of text modalities and motion modalities.

### v5

# A Unified Framework for Motion Reasoning and Generation in Human Interaction

@@ Abstract
Recent advancements in large language models (LLMs) have greatly enhanced their ability to generate natural and contextually relevant text, enabling more human-like AI interactions. However, generating and understanding interactive human-like motion, where multiple individuals engage in coordinated movements, remains challenging due to the complexity of modeling these coordinated interactions. Furthermore, a unified and versatile model is required to handle diverse interactive scenarios, such as chat systems that dynamically adapt to user instructions and assigned roles. To tackle these problems, we introduce VIM, the Versatile Interactive Motion-language model, which integrates both language and motion modalities to effectively understand, generate, and control interactive motions in multi-turn conversational contexts. Unlike previous studies primarily focusing on uni-directional tasks (e.g., text-to-motion or motion-to-text), VIM employs a unified architecture capable of simultaneously understanding and generating both motion and text modalities. Given the lack of an appropriate dataset to address this challenge, we introduce Inter-$\text{MT}^{2}$, a large-scale instruction-tuning dataset containing 82.7K multi-turn interactive motion instructions, spanning 153K interactive motion samples. Inter-$\text{MT}^{2}$ covers diverse instructional scenarios including editing, question answering, and story generation, with interactive motions leveraging off-the-shelf large language models and motion diffusion models. We extensively evaluate the versatility of VIM across multiple interactive motion-related tasks: motion-to-text, text-to-motion, reaction generation, motion editing, and reasoning about motion sequences. Remarkably, VIM is the first model capable of effectively addressing all these tasks with a single unified framework, achieving competitive performance compared to task-specific methods.
Figure 1: We introduce VIM, the Versatile Interactive Motion-language model, a unified architecture that combines language and motion for two-person interactive scenarios. The figure highlights its capabilities across various tasks including motion-to-text, text-to-motion, reaction generation, motion editing, and multi-turn motion reasoning, all within a single framework.
Work done during internship at NAVER AI Lab. Corresponding authors.
@@ 1 Introduction
Modeling interactive human motions stands at the forefront of advancements in robotics and virtual reality. By capturing the subtle nuances of human communications, including gestures, expressions, and interactive behaviors, machines can offer seamless and natural interfaces. This holistic understanding enables technology to adjust its responses and behaviors based on the user’s physical motions and situational context, leading to more personalized and engaging interactions.
@@
Recent advancements in large language models (LLMs) [9, 35, 41] have demonstrated significant potential in generating human-like text and understanding complex linguistic interactions. They have even extended their capability to multi-modal contexts, successfully integrating various input sources such as images, speech, and videos [11, 24, 7, 34, 33]. Building upon these developments, there is a growing interest in incorporating human (or robot) motion as a new modality [19, 6], leading to the emergence of the “motion-language models” (MLM). However, existing approaches [43, 15, 14, 46, 5] often focus on unidirectional tasks that handle one-way translation between text and motion, e.g., text-to-motion or motion-to-text, and consider only single-person motions without interactions. This limitation hinders the agents’ ability to handle scenarios involving interactive motions in multi-turn conversations.
@@
Beyond modeling single-person motions, interactive motions between two individuals allow the model to learn about social behavior. Modeling such interactions requires versatility to effectively control interactions, allowing users to provide instructions, assign roles, or modify behaviors. In this paper, we aim to build a unified yet versatile motion-language model designed to generate, control, and comprehend sophisticated interactive motions.
@@
One of the primary challenges in constructing those models is the lack of multi-turn interactive motion data. Datasets containing motions of two individuals interacting with each other, along with multi-turn conversational instructions, are scarce and challenging to collect. This makes it difficult for models to learn the nuances of interactive motions and multi-turn dynamics. To address this, we present a new interactive motion dataset, Inter-$\text{MT}^{2}$, which contains 82K samples, including various instructional scenarios about the interactive motions in a multi-turn conversational format. We utilize large language models to produce diverse instructions with motion captions and diffusion-based text-to-motion models to generate corresponding interaction motions.
@@
Building upon our Inter-$\text{MT}^{2}$, we present VIM, a Versatile Interactive Motion-language model designed for multi-turn conversations involving interactive motions. We pursue the versatility of VIM through a unified architecture that can simultaneously input and output both motion and text modalities. Based on the pre-trained LLMs, our training process can be divided into three stages: (1) training of the interactive motion tokenizer, (2) pre-training for motion and text representation alignment, and (3) instruction tuning with Inter-$\text{MT}^{2}$ to handle more complex and multi-turn instructions. This enables VIM to effectively comprehend, generate, and control interactive motions, as illustrated in Figure 1. To evaluate VIM’s capabilities, we introduce new protocols that assess its performance on various motion-related tasks, including motion editing and reasoning based on contextual cues, demonstrating its versatility in complex scenarios [^1] We provide more video demos in the supplementary material.. We will publicly release our dataset, codes, and models to facilitate future research.
@@
In summary, the main contributions of this paper are threefold: (1) We propose VIM that can simultaneously process and generate both two-people motion and text modalities, along with a three-stage training pipeline consisting of motion tokenizer training, pre-training for modality alignment, and instruction tuning. (2) We present Inter-$\text{MT}^{2}$, a multi-turn interactive motion-text dataset, to address the lack of multi-turn interactive motion data. (3) We introduce a new evaluation protocol to evaluate the performance of motion-language models on complex motion interaction scenarios.
@@ 2 Related Work
Advancements in human motion modeling have driven significant progress in motion generation and control. Diffusion-based methods [36, 37, 44] have been applied to synthesize human motions from text descriptions. Meanwhile, transformer models using vector quantization [14, 43] have been explored for capturing diverse motion patterns, and MoMASK [15] uses residual tokenizers to enhance fine-grained motion details. For motion editing, some approaches focus on style transfer [1, 16] or specific body part modifications [44, 20]. MEOs [13] use captions and large language models to identify frames and body parts to edit, while MotionFix [2] conditions diffusion models on both source motion and edit text for seamless motion edits. However, these models usually target unidirectional tasks (e.g., text-to-motion, or motion editing) and cannot handle input and output of both motion and text simultaneously in a unified architecture. Unlike existing methods, our approach processes both motion and text concurrently in a unified architecture.
Motion-Language Model
Recent developments in motion-language models have aimed to achieve versatility across various motion-related tasks. MotionGPT [18] demonstrates versatility in motion comprehension and generation based on a unified framework. MotionChain [19] introduces a multi-turn conversational system for interpreting and generating motions within dialogue contexts, including image inputs. Recent work [47, 6, 45, 25] has explored unified approaches to multi-modal motion generation, including speech, video, and image. However, these methods focus on the single-person motions, thus, modeling interactive motions in versatile large models remains under-explored. Wu et al. [38], address the interactive motions, but they still lack multi-turn interactions and complex reasoning abilities. Our work addresses such issues with a model trained on our Inter-$\text{MT}^{2}$ dataset, enabling the understanding and generation of interactive motions in multi-turn conversations with advanced reasoning capabilities.
@@
Modeling human-human interactions has garnered increasing attention in recent research. Several multi-person interaction datasets [26, 10, 42] have been developed, and recent efforts like Inter-X [39] and InterHuman [22] have collected interactive motions paired with textual descriptions for text-based motion control. In text-to-motion tasks [31, 5], InterGEN [39] leverages diffusion with spatial constraint loss, while PriorMDM [32] adapts pre-trained diffusion models with slim communication blocks. For reaction generation, ReMoS [12] uses spatio-temporal cross-attention to synthesize reactive motions, and ReGenNet [40] predicts reactions with a transformer and relative distance-based interaction loss. While existing models have advanced interactive motion modeling, they lack versatility and focus on specific tasks, failing to capture complex multi-turn dynamics. To address this, we introduce Inter-$\text{MT}^{2}$, enabling agents to generate sophisticated motions, respond to instructions, adapt roles, and adjust behaviors based on context.
@@ 3 $\text{Inter}\text{-MT}^{2}$: Interactive multi-turn motion-text dataset
(a) The distribution of instruction types, motion sources, and motion scenario types, highlighting the dataset’s diversity. The type of motion scenario is classified using a large language model with motion captions.
(b) A multi-turn interaction example where two people are playing soccer, illustrating the dataset’s detailed motion and conversational annotations.
Figure 2: Statistics and data sample from Inter-$\text{MT}^{2}$.
@@
In this section, we present Inter-$\text{MT}^{2}$ dataset, for modeling multi-turn interactive motion of multiple humans. Previous datasets [22, 39] provide a textual description of the motions, lack sufficient diversity in instructions, and do not include multi-turn conversations. Since they are insufficient to enable a model to understand and generate complex interaction motions in multi-tern scenarios, we introduce Inter-$\text{MT}^{2}$: Interactive Muti-Turn Motion-Text dataset. This dataset covers a variety of interactive motion scenarios with multi-turn conversations, diverse instructions, and spatiotemporally aligned motions between two individuals. We enhance our dataset by generating diverse instructions from large language models and combining motion data from existing datasets with generative approaches to enable flexible text-to-motion modeling.
Dataset Ret. top-3 Div.
Source dataset 0.870 0.997
Generated by InterGEN 0.645 0.953
Inter-$\text{MT}^{2}$ (Ours) 0.701 0.931
Table 1: Comparison of generated motions on text-matching ability (top-3 retrieval precision), and motion diversity (Div.).
We begin with the human interaction motion and text datasets, Inter-X [39] and InterHuman [22], as the foundational resources for our dataset construction. To convert these datasets into instructional datasets, we first generate multi-turn instructions with motion captions using GPT-4o [27]. We consider the instructional scenarios as various tasks with following text prompts, including motion editing (e.g., “Make the left person more playful”), motion reasoning (e.g., “What happened before/after this motion?”), and story generation (e.g., “Let’s create a story where two people are following this motion.”). Detailed prompt templates and the complete data collection pipeline are presented in the supplementary materials. To guarantee high-quality caption generation, we guide the LLMs by providing action labels from the existing datasets alongside example captions, effectively constraining and enhancing the relevance and accuracy of the generated captions. Subsequently, we utilize a state-of-the-art diffusion-based text-to-motion model, InterGEN [22], to synthesize interactive motions that align closely with these generated captions.
Our pipeline creates samples in two ways. First, starting with a dataset motion, we generate a caption and instruction and then use InterGEN [22] to synthesize a matching motion, yielding both the original and synthesized motions with the instruction. Alternatively, we generate two captions and instructions to synthesize two motions, producing samples entirely from synthesized motions. This method blends data-sourced and generative motions for reliable interactive motion modeling. Overall, we collected 82K multi-turn conversations, including 96K synthesized and 56K real motions. Figure 2 shows statistics and samples from our Inter-$\text{MT}^{2}$, where motion scenarios are classified using a large language model with motion captions.
To assess the quality and diversity of the generated motions and their alignment with texts, we evaluate our dataset using the text-motion matching score and diversity metric of our dataset, as shown in Table 1. Pre-trained retrieval models [29] assess the alignment between motions and captions, with additional details in the supplementary material. Our dataset achieves a top-3 retrieval precision of $0.701$ (the precision of the source dataset of the retrieval model is $0.870$), showing good alignment, which slightly surpasses the matching performance of the synthesized dataset created by the state-of-the-art motion generation method, InterGEN [22]. Additionally, our dataset exhibits robust diversity similar to the source dataset. These results indicate that despite our multi-turn interactive motions and captions being synthetically generated, their quality closely approximates that of real-world datasets.
Figure 3: An overview of VIM, illustrating its versatile capability to flexibly process and generate interactive motions and texts in an auto-regressive manner. We omit the motion tokenizer, which converts raw motion sequences into discrete motion tokens, for clarity. VIM covers versatile motion tasks involving both motion and textual modalities across multiple conversational turns.
@@ 4 VIM: Versatile Interactive Motion-Language Model
In this section, we introduce VIM, a versatile interactive motion-language model that processes multi-turn conversations with both language and two-person interactive motions as inputs and outputs. First, we will explain our design choices for the model architectures, followed by a detailed description of the training methodologies.
@@ 4.1 Notations
We denote an interactive motion from two individual $a$ and $b$ as $\{\mathbf{m}_{a},\mathbf{m}_{b}\}$, following non-canonical representation in [22] based on SMPL-X structure [28] with $M$ as a motion length. At each motion time step $i$, the motion representation is defined as: $\mathbf{m}^{i}=[\mathbf{j}_{g}^{p},\mathbf{j}_{g}^{v},\mathbf{j}^{r},\mathbf{c}^{f}]$, where $\mathbf{j}_{g}^{p}\in\mathbb{R}^{3N_{j}}$ is the global joint positions, $\mathbf{j}_{g}^{v}\in\mathbb{R}^{3N_{j}}$ is the global joint velocities, $\mathbf{j}^{r}\in\mathbb{R}^{6N_{j}}$ is 6D representation of local rotations with $N_{j}$ joints, and $\mathbf{c}^{f}\in\mathbb{R}^{4}$ is binary ground contact features. We train a motion-language model $p_{\theta}$ that jointly models text and motion data. The model processes the input (user instructions or context) and output (machine responses), effectively integrating both modalities.